# Patterns vs. Patients: Evaluating LLMs against Mental Health Professionals on Personality Disorder Diagnosis through First-Person Narratives

### v1

@@ Abstract
Growing reliance on LLMs for psychiatric self-assessment raises questions about their ability to interpret qualitative patient narratives. We present the first direct comparison between state-of-the-art LLMs and mental health professionals in diagnosing Borderline (BPD) and Narcissistic (NPD) Personality Disorders utilizing Polish-language first-person autobiographical accounts. We show that the top-performing Gemini Pro models surpassed human professionals in overall diagnostic accuracy by 21.91 percentage points (65.48% vs. 43.57%). While both models and human experts excelled at identifying BPD ($F_{1}=83.4$ & $F_{1}=80.0$, respectively), models severely underdiagnosed NPD ($F_{1}=6.7$ vs. $50.0$), showing a reluctance toward the value-laden term ,,narcissism.” Qualitatively, models provided confident, elaborate justifications focused on patterns and formal categories, while human experts remained concise and cautious, emphasizing the patients’ sense of self and temporal experience. Our findings demonstrate that while LLMs are highly competent at interpreting complex first-person clinical data, their outputs still carry critical reliability and bias issues.
@@ 1 Introduction
Systemic pressures on mental healthcare have created critical accessibility gaps and diagnostic delays Sun2023; Barbui2025. Consequently, the public is increasingly turning to widely available Large Language Models (LLMs) for self-assessment, thereby bypassing traditional clinical pathways Lawrence2024; McBain2025. While this trend may partially relieve strain on the healthcare system, it also introduces considerable ethical and clinical risks, such as the potential for inaccurate or misleading diagnoses Harrer2023. This study empirically compares the diagnostic capabilities of current open- and closed-source LLMs to mental health professionals in identifying a complex and frequently misdiagnosed condition of Personality Disorder (PD).
@@
Previous research shows that state-of-the-art LLMs can achieve performance levels that approach, or in some cases surpass, those of trained clinicians on standardized medical benchmarks grzybowski-etal-2025-polish; workum2025, and substantially improve clinicians’ diagnostic accuracy when used as a decision-support tool noda2025. These capabilities extend to chatbot-patient interaction, where recent investigations suggest that LLMs can surpass human physicians in expressing empathy in written clinical communication ayers2023. However, it remains unclear whether such capabilities can be effectively applied to complex psychiatric conditions, where diagnoses often lack well-established objective markers and depend heavily on nuanced, subjective accounts of patients.
@@
A limited body of work evaluating LLMs specifically on psychiatric and psychological benchmarks reveals a more uneven performance. Studies show that while models such as GPT-4 and Llama 3 achieve high accuracy (up to 85%) in binary mental disorder detection tasks, their performance varies significantly across different datasets hanafi2025comprehensiveevaluationlargelanguage. Notably, LLMs fine-tuned on domain-specific data do not outperform generalist models on existing benchmarks fouda2025psychiatrybenchmultitaskbenchmarkllms; nguyen-etal-2025-large.
Despite these advances, most existing evaluations rely on binary or multiple-choice answer formats and use clinical-annotated datasets or noisy social media posts. Such evaluations are not directly comparable to the semi-structured, first-person narrative through which patients typically communicate their experiences to a mental health professional. Furthermore, existing studies have prioritized outcome metrics over explanatory evaluations of the reasoning process. Consequently, they do not verify whether models emulate human cognitive processes or if their justifications correspond to human diagnostic decisions. This discrepancy highlights a critical research gap: it remains unknown if LLMs can maintain diagnostic accuracy on raw first-person testimonies, provide diagnosis-relevant justifications, and whether their performance and reasoning compare with those of mental health professionals.
We explore this gap in a specific context of two conditions, namely Borderline Personality Disorder (BPD) and Narcissistic Personality Disorder (NPD). These are strong, yet underexplored, cases for model evaluation, particularly as their diagnostic assessment might be overdetermined by the theoretical framework applied, or the lack thereof (as with the lay understanding of the terms ,,borderline” and ,,narcissism”).
While the DSM-III  APA1980 formally established Personality Disorder (PD) as a standalone diagnostic unit, its initial categorical approach—in which mental health disorders are perceived as binary categories—has been criticized for producing high comorbidity and permitting substantial heterogeneity within categories Widiger2007. In response, recent revisions in DSM-5 APA2022 and ICD-11 WHO2022 have redirected attention from distinct categories of PD to transdiagnostic commonalities and their continuous variation, i.e., a dimensional, non-exclusive approach. The DSM-5 AMPD (Section III) APA2022 introduced a general criterion (Criterion A) that defines a shared basis for all PD, independent of their categorical manifestations Pincus2020.
@@
Given the prominence of the ICD and DSM frameworks, current LLMs could be biased toward a dimensional understanding of PD. Conversely, the ubiquity of terms like ,,borderline” and ,,narcissism” in everyday language—and thus in training corpora—suggests a potential counter-bias toward a categorical understanding of PD. Since empirical evidence does not clarify whether current LLMs lean toward a dimensional or categorical view of PD, we examined both frameworks independently.
@@
  • We adopt an explanatory approach to both human data and model outputs, aiming not only to quantify performance but also to provide interpretable insights.
@@ 2.1 Data
To ensure the clinical validity of the sensitive first-person data used for evaluation, all narratives were collected at the Psychiatric Hospital in Międzyrzecz, Poland, following a multi-stage procedure.
First, 120 patients with a diagnosed personality disorder were identified. Each diagnosis had been established in outpatient care and subsequently confirmed by two experienced psychiatrists during inpatient admission, followed by an assessment by a clinical psychologist. Second, cases involving suicidal crisis, psychotic decompensation, or comorbid substance use disorders were excluded. Third, we employed intensity sampling to select information-rich, yet non-extreme, cases suitable for both qualitative and computational analysis.
Based on ICD-10 diagnostic coding, 24 BPD and 20 NPD cases were shortlisted. To ensure comparability, we assessed personality functioning and maladaptive traits using a multi-dimensional protocol (see Appendix A). All patients participated in 50–70 minute semi-structured qualitative interviews adapted from McAdams’ Life Story Interview McAdams1988, designed to explore narrative identity. Two additional questions probing reflective self-experience were included. Interviews were audio-recorded and transcribed, resulting in a Polish-language corpus exceeding 200,000 words. The qualitative analyses of all BPD cases Sterna2025 and all NPD cases (conducted for this project) provided an in-depth understanding of the subjective nuances represented in the data.
Six cases—three BPD and three NPD—were selected for this study. These were meticulously chosen to represent mild, moderate, and severe levels of impairment, and to capture the breadth of narrative expression within each disorder. Their selection reflects both the unique richness of their narratives and the practical constraints imposed by a human-model comparison (since each narrative is long, complex, and requires significant time commitment and detailed evaluation).
Additionally, from a pool of 20 control participants who underwent the same interview procedure, one control case with no personality disorder and subclinical personality functioning was chosen. In total, the evaluation dataset comprised 7 transcriptions of autobiographical interviews: 3 BPD, 3 NPD, and 1 Healthy Control.
@@ 2.2 Model Selection
We included a total of N=16{N=16} leading models and variants, split into a closed-source ($N=9$) and an open-source group ($N=7$). Our selection criteria encompassed: general public interest and availability, presence in recent evaluations on related material fouda2025psychiatrybenchmultitaskbenchmarkllms; hua2025scoping; nguyen-etal-2025-large, access model (licensing), parameter count, country of origin, and reasoning ability.
The closed-source group included: Gemini 2.5 Pro comanici2025gemini25pushingfrontier, Gemini 3 Pro google2025gemini3, Claude Opus 4.1 anthropic2025claude4, GPT-4o openai2024gpt4ocard, GPT-4.1 openai2025gpt41, and four variants of GPT-5 openai2025gpt5 with all available values of the OpenAI API’s reasoning effort parameter (from minimal to high). The open-source group consisted of: Gemma 3 27B gemmateam2025gemma3technicalreport, Llama 3.3 70B grattafiori2024llama3herdmodels, DeepSeek R1 0528 deepseekai2025deepseekr1incentivizingreasoningcapability, two variants of DeepSeek v3.1 Terminus deepseekai2025deepseekv3technicalreport with reasoning either enabled or disabled, and two variants of Qwen 3 32B yang2025qwen3technicalreport, also with and without reasoning.
We did not include domain-specific models trained on medical data for two reasons. Firstly, the context windows of available medical models evaluated in existing literature fouda2025psychiatrybenchmultitaskbenchmarkllms; hanafi2025comprehensiveevaluationlargelanguage; hua2025scoping; nguyen-etal-2025-large proved to be too small for the full patient testimonies—except for MentalQLM shi2025mentalqlm, which has not been made publicly available. Secondly, recent studies indicate that general-purpose models perform similarly to or better than medical models on psychiatric tasks fouda2025psychiatrybenchmultitaskbenchmarkllms.
@@ 2.3 Mental Health Experts
The study sample consisted of $N=6$ highly experienced mental health professionals, comprising three psychiatrists and three psychotherapists, recruited through our professional network. The group was balanced by gender, with three participants identifying as men and three as women; the mean age was 48.5 years (SD = 11.18), with an average of 18.5 years of professional experience. Importantly, the experts were external to the original diagnostic teams. They were blinded to the patients’ previous medical records and ground truth diagnoses, and had never previously encountered the patients whose narratives they evaluated.
@@ 2.5.2 Diagnostic Justifications
Human mental health professionals can arrive at similar diagnostic conclusions through different cognitive and emotional pathways biondi2022clinician. Traces of these mental processes are reflected in the semantic content of the justifications provided during the diagnosis. To use this information alongside quantitative metrics, we mapped the diagnostic justifications onto a high-dimensional semantic embedding space. This allowed for a comparison of diagnostic expertise leverage between human and model participants, as well as among different models.
BAAI/bge-multilingual-gemma2111https://huggingface.co/BAAI/bge-multilingual-gemma2 was chosen as the embedding model owing to its superior performance on Polish-language tasks in the MMTEB benchmark enevoldsen2025mmtebmassivemultilingualtext. We created a single summary embedding representing the averaged semantic content of justifications for each model separately, as well as a single one representing all human participant data to account for its comparative scarcity. A more detailed process of embedding creation was described in Appendix G. The summary embeddings were aggregated into two distinct datasets: a global dataset comprising all agents (both models and humans) to assess the human-AI semantic gap, and a model-only dataset to allow for a finer-grained analysis of inter-model semantic differences.
To visualize the high-dimensional relationships between these representations, we employed a dual-method dimensionality reduction approach utilizing Multidimensional Scaling (MDS) kruskal1964multidimensional and Uniform Manifold Approximation and Projection (UMAP) mcinnes2020umapuniformmanifoldapproximation. This combination was chosen to provide a comprehensive view of the semantic space: MDS was selected for its ability to preserve global metric distances, offering a faithful representation of the relative dissimilarities between models. Complementarily, UMAP was utilized for its strength in manifold learning and local structure preservation, which allows for the identification of distinct semantic clusters (neighborhoods).
MDS was conducted using the scikit-learn [^2] https://github.com/scikit-learn/scikit-learn library JMLR:v12:pedregosa11a. Before dimensionality reduction, a pairwise cosine distance matrix was computed for the embeddings to ensure that the angular relationships inherent to semantic vector spaces were preserved. UMAP was conducted using the umap-learn [^3] https://github.com/lmcinnes/umap library. Details on the hyperparameter values used can be found in Appendix F.
@@ 2.5.3 Lexical Features
To further investigate the divergence between linguistic justification of diagnostic reasoning in humans and models, we conducted a follow-up inquiry into interpretable differences. We aggregated all justifications into two distinct corpora—human-written and model-written—and sought to identify lexical features that were statistically overrepresented in one group relative to the other. The features were identified with the weighted log-odds ratio with informative Dirichlet prior method originally applied to the problem of detecting lexical polarization in political discourse monroe2008fightin.
In the technical implementation, we first applied a text pre-processing pipeline to the two corpora, converting non-Polish non-ASCII characters into ASCII equivalents, as well as removing LLM-characteristic text formatting artifacts, punctuation, digits, and optionally stop words. Finally, we applied Monroe et al.’s method through the ConvoKit library chang-etal-2020-convokit. We considered two parameters to adjust in the feature generation process: stop word removal—either enabled or disabled, and the maximum n-gram length—exclusive ranges of 1 to 3 and a combined 1–3 one, ultimately choosing unigrams with the stop words removed as the most informative configuration.
@@ 3.1 Diagnostic Performance
Best-performing models outperform the human diagnostic average. The overall human average score was 43.57%, with a categorical score of 3.60 out of 7 (51.43%) and a dimensional score of 2.50 out of 7 (35.71%). An exploratory analysis of the participant sub-groups suggested a potential performance gap: psychotherapists scored higher on the categorical task (mean difference: 2.33), while psychiatrists scored slightly higher on the dimensional task (mean difference: 0.33).
@@
Further analysis of model performance indicated that reasoning did not consistently improve diagnostic accuracy. Model size appeared to be a factor, with smaller models generally performing worse than larger ones. No clear performance gap was observed between closed-source and open-source models.
BPD overdiagnosis and NPD underdiagnosis drive systematic bias, most pronounced in models. An analysis of the $N=378$ total diagnoses (see Appendix C, Table C.1) by both models ($N=336$) and humans ($N=42$) revealed systematic diagnostic bias.
@@
The ,,healthy” label was the second most common, accounting for 21.43% of all assigned diagnoses. Models were highly effective at identifying the absence of pathology ($recall=87.5\%$, $F1=69.4$) compared to human experts ($recall=60.0\%$, $F1=50.0$). Despite these high scores, a detailed analysis of diagnostic distributions (see Appendix D, Figure D.1) reveals a specific ,,depathologizing bias” in the GPT family. GPT-4o showed the most extreme bias, misclassifying cases as healthy 10 times. GPT-4.1 followed with 5 false positives, whereas the GPT-5 variants had 2–3 each. This suggests a tendency in the GPT model family to favor non-clinical classifications. Notably, humans also showed a slight bias toward non-clinical classifications, evidenced by 4 false positives (see Appendix D, Figure D.1), driven by frequent assessments that the testimonies lacked sufficient evidence for a PD diagnosis.
A significant misclassification bias was observed for Avoidant Personality Disorder (AvPD) (see Appendix C, Table C.1). It emerged as the third most common diagnosis overall (15.08%), but this prevalence was driven largely by models. AvPD accounted for 16.07% (54 of 336) of model-generated diagnoses, compared to 7.14% (3 of 42) of human diagnoses. Given that AvPD was not present in the ground truth answers, this high prevalence points to a systematic tendency to misinterpret symptoms of other conditions as AvPD. On the other hand, only one out of 336 diagnoses by models indicated a PD category not present in current diagnostic frameworks, namely Masochistic PD.

### v2

@@ Abstract
Growing reliance on LLMs for psychiatric self-assessment raises questions about their ability to interpret qualitative patient narratives. This depth-first case study provides the first direct comparison of state-of-the-art LLMs and mental health professionals in assessing Borderline (BPD) and Narcissistic (NPD) Personality Disorders based on Polish-language first-person autobiographical accounts. Within our sample, the overall diagnostic scores of the top-performing Gemini Pro models (65.48%) were 21.91 percentage points higher than the average scores of the human professionals (43.57%). While both models and human experts excelled at identifying BPD ($F_{1}=83.4$ & $F_{1}=80.0$, respectively), models severely underdiagnosed NPD ($F_{1}=6.7$ vs. $50.0$), showing a potential reluctance toward the value-laden term ,,narcissism.” Qualitatively, models provided confident, elaborate justifications focused on patterns and formal categories, while human experts remained concise and cautious, emphasizing the patients’ sense of self and temporal experience. Our findings demonstrate that while LLMs might be competent at interpreting complex first-person clinical data, their outputs still carry critical reliability and bias issues.
@@ 1 Introduction
Systemic pressures on mental healthcare have led to critical access gaps and diagnostic delays Sun et al. (2023); Barbui et al. (2025). Consequently, the public is increasingly turning to widely available Large Language Models (LLMs) for self-assessment, thereby bypassing traditional clinical pathways Lawrence et al. (2024); McBain et al. (2025). While this trend may partially relieve strain on the healthcare system, it also introduces considerable ethical and clinical risks, such as the potential for inaccurate or misleading diagnoses Harrer (2023). This study empirically compares the diagnostic capabilities of current open- and closed-source LLMs to mental health professionals in identifying a complex and frequently misdiagnosed condition of Personality Disorder (PD).
@@
Previous research shows that state-of-the-art LLMs can achieve performance levels that approach, or in some cases surpass, those of trained clinicians on standardized medical benchmarks Grzybowski et al. (2025); Workum et al. (2025), and substantially improve clinicians’ diagnostic accuracy when used as a decision-support tool Noda et al. (2025). These capabilities extend to chatbot-patient interaction, where recent investigations suggest that LLMs can surpass human physicians in expressing empathy in written clinical communication Ayers et al. (2023). However, it remains unclear whether such capabilities can be effectively applied to complex psychiatric conditions, where diagnoses often lack well-established objective markers and depend heavily on nuanced, subjective accounts of patients.
@@
A limited body of work evaluating LLMs specifically on psychiatric and psychological benchmarks reveals a more uneven performance. Studies show that models such as GPT-4 and Llama 3 achieve high accuracy (up to 85%) on binary mental disorder detection tasks, but their performance varies significantly across datasets Hanafi et al. (2025). Notably, LLMs fine-tuned on domain-specific data do not outperform generalist models on existing benchmarks Fouda et al. (2025); Nguyen et al. (2025).
Existing evaluations typically rely on simplified formats and noisy or synthetic datasets, which are not comparable to the complex, first-person nature of psychiatric patients’ testimonies. Furthermore, by prioritizing outcome metrics over explanatory reasoning, current studies fail to verify whether models emulate human cognitive processes or merely guess accurately. To address this gap, our study introduces a methodological shift from breadth to depth. Specifically, instead of evaluating large volumes of surface-level material, we deliberately focus on a smaller number of raw, in-depth clinical cases. We apply this depth-first approach to Borderline Personality Disorder (BPD) and Narcissistic Personality Disorder (NPD). These are strong, yet underexplored, phenomenons for model evaluation, particularly as their diagnostic assessment might be overdetermined by the theoretical framework applied (or the lack thereof, as with the lay understanding of the terms ,,borderline” and ,,narcissism”).
@@
Recent revisions in DSM-5 American Psychiatric Association (2022) and ICD-11 World Health Organization (2022) have redirected attention from a distinct categorical approach to PD to one based on transdiagnostic commonalities and their continuous variation, i.e., a dimensional, non-exclusive approach. Given the prominence of the ICD and DSM frameworks, current LLMs could be biased toward a dimensional understanding of PD. Conversely, the ubiquity of terms like ,,borderline” and ,,narcissism” in everyday language—and thus in training corpora—suggests a potential counter-bias toward a categorical understanding of PD. Since empirical evidence does not clarify whether current LLMs lean toward a dimensional or categorical view of PD, we examined both frameworks independently.
@@
  • We prioritize depth over breadth, adopting an explanatory approach to both human data and model outputs, aiming not only to quantify performance but also to provide interpretable insights.
@@ 2.1 Data
The database comprises six clinical cases—three BPD and three NPD—alongside one healthy control. These were selected from a larger dataset of first-person symptomatology, which includes 24 BPD cases, 20 NPD cases, and 20 healthy controls (HC), classified according to ICD-10 diagnostic criteria.
To ensure the clinical validity of the sensitive first-person data used for evaluation, clinical narratives were collected at the Psychiatric Hospital in Międzyrzecz, Poland, following a multi-stage procedure. Participants enrolled in the original study were diagnosed as outpatients, then confirmed as inpatients; each diagnosis was confirmed by four independent clinical experts: two psychiatrists and two psychologists. To ensure comparability, categorical diagnosis was accompanied by personality functioning and maladaptive traits assessment, using a multi-dimensional protocol (see Appendix A). HC participants were healthy volunteers who responded to the study’s announcement and declared no diagnosed psychiatric conditions or felt symptoms. They were subsequently screened for personality pathology using LPFS BF 2.0  Łakuta et al. (2023) and PICD Cieciuch et al. (2022).
All the participants took part in 50–70 minute semi-structured qualitative interviews adapted from McAdams’ Life Story Interview McAdams (1988), designed to explore narrative identity. Two additional questions probing reflective self-experience were included. Interviews were audio-recorded and transcribed, resulting in a Polish-language corpus exceeding 200,000 words.
To extract cases for this study, we followed intensity sampling to select information-rich, yet non-extreme cases representing: integrated personality as defined in DSM-5 (HC, $N=1$) and mild, moderate, and severe levels of personality impairment (clinical sample: BPD, $N=3$; NPD, $N=3$). The severity variance represented in the sample was meant to capture the breadth of narrative expression within both BPD and NPD.
The final sample size was set to balance the necessity to represent the heterogeneity of PD symptomatology with the practical constraints imposed by a human evaluation. The final database, amounting to 27,641 words, was evaluated for approximately five hours per case on average, thus reaching the maximum volume that avoided overwhelming human evaluators and secured the reliability of their assessments.
@@ 2.2 Model Selection
We included a total of N=16{N=16} leading models and variants, split into a closed-source ($N=9$) and an open-source group ($N=7$). Our selection criteria encompassed: public interest and availability, presence in recent evaluations on related material Fouda et al. (2025); Hua et al. (2025); Nguyen et al. (2025), access model (licensing), parameter count, country of origin, and reasoning ability.
The closed-source group included: Gemini 2.5 Pro Gemini Team (2025), Gemini 3 Pro Google (2025), Claude Opus 4.1 Anthropic (2025), GPT-4o OpenAI (2024), GPT-4.1 OpenAI (2025b), and four variants of GPT-5 OpenAI (2025a) with all available values of the OpenAI API’s reasoning effort parameter (from minimal to high). The open-source group consisted of: Gemma 3 27B Gemma Team (2025), Llama 3.3 70B Llama Team (2024), DeepSeek R1 0528 DeepSeek-AI (2025a), two variants of DeepSeek v3.1 Terminus DeepSeek-AI (2025b) with reasoning either enabled or disabled, and two variants of Qwen 3 32B Qwen Team (2025), also with and without reasoning.
We did not include domain-specific models trained on medical data for two reasons. Firstly, the context windows of available medical models evaluated in existing literature Fouda et al. (2025); Hanafi et al. (2025); Hua et al. (2025); Nguyen et al. (2025) proved to be too small for the full patient testimonies—except for MentalQLM Shi et al. (2025), which has not been made publicly available. Secondly, recent studies indicate that general-purpose models perform similarly to or better than medical models on psychiatric tasks Fouda et al. (2025).
@@ 2.3 Mental Health Experts
The study sample consisted of $N=6$ highly experienced mental health professionals, comprising three psychiatrists and three psychotherapists, recruited through our professional network. They received no compensation for their participation in the study. The group was balanced by gender, with three participants identifying as men and three as women; the mean age was 48.5 years (SD = 11.18), with an average of 18.5 years of professional experience. Importantly, the experts were external to the original diagnostic teams. They were blinded to the patients’ previous medical records and ground truth diagnoses, and had never previously encountered the patients whose narratives they evaluated. Each expert formulated their diagnostic decisions entirely independently.
@@ 2.4 Procedure
Importantly, to reliably evaluate the baseline competence of the models, no model-only in-context learning paradigm was employed in the protocol. This is because potential differences in understanding the patients’ data between human participants and models were another point of interest of the evaluation. The differences could potentially stem from, e.g., a lack of such priming through example human responses.
@@ 2.5.2 Diagnostic Justifications
Human mental health professionals can arrive at similar diagnostic conclusions through different cognitive and emotional pathways Biondi et al. (2022). Traces of these mental processes are reflected in the semantic content of the justifications provided during the diagnosis. To use this information alongside quantitative metrics, we mapped the diagnostic justifications onto a high-dimensional semantic embedding space. This allowed for a comparison of diagnostic expertise leverage between human and model participants, as well as among different models.
BAAI/bge-multilingual-gemma2111https://huggingface.co/BAAI/bge-multilingual-gemma2 was chosen as the embedding model owing to its superior performance on Polish-language tasks in the MMTEB benchmark Enevoldsen et al. (2025). We created a single summary embedding representing the averaged semantic content of justifications for each model separately, as well as a single one representing all human participant data to account for its comparative scarcity. A more detailed process of embedding creation was described in Appendix G. The summary embeddings were aggregated into two distinct datasets: a global dataset comprising all agents (both models and humans) to assess the human-AI semantic gap, and a model-only dataset to allow for a finer-grained analysis of inter-model semantic differences.
Due to the depth-first approach of the study, the summary embeddings constituted an unusually small set of 17 data points, making a faithful visualization of the high-dimensional relationships between these representations harder. Instead of the domain standard methods such as UMAP McInnes et al. (2020), t-SNE van der Maaten and Hinton (2008), or PCA Pearson (1901), we employed Multidimensional Scaling (MDS) Kruskal (1964) as the dimensionality reduction algorithm. Unlike local manifold techniques—unreliable for small sample sizes—or linear projections—which prioritize variance over similarity—MDS was selected for its ability to preserve global metric distances, offering a faithful representation of the relative semantic dissimilarities between the summary embeddings. The algorithm was implemented using the scikit-learn [^2] https://github.com/scikit-learn/scikit-learn library Pedregosa et al. (2011). Before dimensionality reduction, a pairwise cosine distance matrix was computed for the embeddings to ensure that the angular relationships inherent to semantic vector spaces were preserved. Details on the hyperparameter values used can be found in Appendix F.
@@ 2.5.3 Lexical Features
To further investigate the divergence between linguistic justification of diagnostic reasoning in humans and models, we conducted a follow-up inquiry into interpretable differences. We aggregated all justifications into two distinct corpora—human-written and model-written—and sought to identify lexical features that were statistically overrepresented in one group relative to the other. The features were identified with the weighted log-odds ratio with the informative Dirichlet prior method originally applied to the problem of detecting lexical polarization in political discourse Monroe et al. (2008). The method is well-suited for analyzing small and/or imbalanced text corpora; it estimates the z-score of the log-odds ratio for each n-gram, controlling for variance in word frequency, which specifically addresses the issue of sample size imbalance between the human and model datasets in our use case.
In the technical implementation, we first applied a text pre-processing pipeline to the two corpora, converting non-Polish non-ASCII characters into ASCII equivalents, as well as removing LLM-characteristic text formatting artifacts, punctuation, digits, and optionally stop words. Finally, we applied Monroe et al.’s method through the ConvoKit library Chang et al. (2020). We considered two parameters to adjust in the feature generation process: stop word removal—either enabled or disabled, and the maximum n-gram length—exclusive ranges of 1 to 3 and a combined 1–3 one, ultimately choosing unigrams with the stop words removed as the most informative configuration.
@@ 3.1 Diagnostic Performance
Top-performing models achieve a higher diagnostic average than human experts. The overall human average score was 43.57%, with a categorical score of 3.60 out of 7 (51.43%) and a dimensional score of 2.50 out of 7 (35.71%). An exploratory analysis of the participant sub-groups suggested a potential performance gap: psychotherapists scored higher on the categorical task (mean difference: 2.33), while psychiatrists scored slightly higher on the dimensional task (mean difference: 0.33). Importantly, inter-rater reliability among human experts was relatively low, reflecting the inherent subjectivity of psychiatric evaluation Di Forti et al. (2025) restricted to the text modality Fiquer et al. (2013). For categorical diagnosis, Fleiss’ Kappa was 0.11, and Krippendorff’s Alpha (nominal) was 0.14, indicating slight agreement; for dimensional severity ratings, Krippendorff’s Alpha (ordinal) was 0.22, indicating fair agreement.
@@
Further analysis of model performance indicated that reasoning did not consistently improve diagnostic accuracy. Additionally, no clear performance gap was observed between closed-source and open-source models. In contrast, model size appeared to be a factor, with smaller models generally performing worse than larger ones. However, drawing definitive statements about its influence is difficult, as they would rely exclusively on the open-sourced models in our sample. This is further compounded by the fact that model performance being inevitably correlated with model size is (mostly) a general principle of machine learning.
Diagnostic distributions reveal frequent BPD overdiagnosis and NPD underdiagnosis tendencies, particularly among models. An analysis of the $N=378$ total diagnoses (see Appendix C, Table C.1) by both models ($N=336$) and humans ($N=42$) revealed systematic diagnostic bias.
@@
The ,,healthy” label was the second most common, accounting for 21.43% of all assigned diagnoses. Models were highly effective at identifying the absence of pathology ($recall=87.5\%$, $F1=69.4$) compared to human experts ($recall=60.0\%$, $F1=50.0$). Despite these high scores, a detailed analysis of diagnostic distributions (see Appendix D, Figure D.1) reveals a specific ,,depathologizing bias” in the GPT family. GPT-4o showed the most extreme bias, misclassifying cases as healthy 10 times. GPT-4.1 followed with 5 false positives, whereas the GPT-5 variants had 2–3 each. This suggests a tendency in the GPT model family to favor non-clinical classifications. Notably, humans also demonstrated a slight tendency toward non-clinical classifications, evidenced by 4 false positives (see Appendix D, Figure D.1), driven by frequent assessments that the testimonies lacked sufficient evidence for a PD diagnosis.
A significant misclassification bias was observed for Avoidant Personality Disorder (AvPD) (see Appendix C, Table C.1). It emerged as the third most common diagnosis overall (15.08%), but this prevalence was driven largely by models. AvPD accounted for 16.07% (54 of 336) of model-generated diagnoses, compared to 7.14% (3 of 42) of human diagnoses. Given that AvPD was not present in the ground truth answers, this high prevalence points to a tendency to misinterpret symptoms of other conditions as AvPD. On the other hand, only one out of 336 diagnoses by models indicated a PD category not present in current diagnostic frameworks, namely Masochistic PD.